# A shared-revenue Bertrand game

### v1

@@ 3.4. Summary of Nash equilibria
Equilibrium Price Configurations of Shared-Revenue Bertrand Game
@@ 4.2. Relative Pareto optimality
$(p_{r},p_{s})\in\left(p_{r^{*}},\left[p^{\dagger},p_{s^{*}}\right]\right)\cup\left(p_{s},\left[p_{s_{\text{ind}}},p^{\dagger}\right]\right)\ ;$ if $\text{if }\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\text{ and }p^{\dagger}\leq p_{s^{*}}$
The equilibrium where the retailer fulfills demand at $p_{r^{*}}$ is Pareto suboptimal relative to the shared-price equilibrium $p^{\dagger}$ where the independent seller fulfills demand. Indeed, by construction of $p^{\dagger}$, in both cases the retailer achieves the same payoff of $\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$ so they are indifferent. However, because $\alpha\leq\alpha^{\dagger}_{s}$ implies that $p_{s_{\text{ind}}}\leq p^{\dagger}$, the independent seller achieves a positive payoff when fulfilling demand at $p^{\dagger}$ while they achieve a payoff of 00 when allowing the retailer to fulfill demand at $p_{r^{*}}$. Thus, they strictly prefer fulfilling demand at $p^{\dagger}$ rather than allowing the retailer to fulfill demand at $p_{r^{*}}$.
As a remark, we note that a reasonable argument could be made for preferring the equilibrium where the retailer fulfills demand at $p_{r^{*}}$. In this case $p_{r^{*}}<p^{\dagger}$, so despite the Pareto optimality argument, the customer will be able to buy the product at a lower price if the retailer fulfills demand at $p_{r^{*}}$ instead of the independent seller fulfilling demand at $p^{\dagger}$. With this lowest price criterion, $p_{r^{*}}$ would in fact be the most likely equilibrium in this case, as $p_{r^{*}}<p_{s_{\text{ind}}}$ as well. However, whatever we choose does not substantially affect the results, and the relative Pareto optimality criteria will continue to be sensible for us as we generalize the game, thus we proceed with this assumption.
@@ 4.3. Equilibrium outcomes
Equilibrium Outcomes of Shared-Revenue Bertrand Game
@@
A quick glance at (26) and (27) shows that when $\alpha$ is sufficiently high, the independent seller is “priced out” of the market and the retailer will fulfill demand at their single agent price as if the independent seller did not exist. When $\alpha$ is small, the independent seller sells at the retailer’s indifference price, illustrating that they have to prioritize preventing the retailer from undercutting them over maximizing their own payoff. However, when $\alpha$ is in a certain sweet spot, the independent seller may be able to sell at their optimal price in equilibrium. We will see soon that everybody wins in this regime – if the retailer gets to choose the referral fee, they would prefer to choose a referral fee in this sweet spot rather than one that is too high or too low.
@@ 4.4. Fee and cost structure analysis
4.4.1. $(p_{r^{*}},r)$ The retailer fulfills demand at its single agent price
@@
4.4.3. $(p_{r_{\text{ind}}},s)$ The independent seller fulfills demand at the retailer’s indifference price, $p_{r_{\text{ind}}}$
@@
4.4.4. $(p_{s^{*}},s)$ The independent seller fulfills demand at their single agent price, $p_{s^{*}}$
When $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$, we have an equilibrium in which the independent seller fulfills demand at its optimal price $p_{s^{*}}$ (red region). In this regime, we always have $p_{s^{*}}\leq p_{r^{*}}$, thus the customers are able to purchase at a lower price than they would be able to without the revenue sharing program. Of course, the independent seller prefers this equilibrium because the fee is fairly low and they are able to sell at their optimal price. Finally, it’s even possible that there exists some $\alpha$ in the red region such that the retailer achieves a payoff higher than their optimal single agent payoff, yielding a win-win-win situation for the customers, independent seller, and retailer.
@@ 5. Optimizing the referral fee
In Section 4.4, we identified an equilibrium that is a win-win-win (red region) for the customers, independent seller, and retailer, along with a “safe” equilibrium (gray region) for the retailer to achieve a payoff no less than its single agent payoff. While we have made an intuitive argument about which equilibria are preferable, the next step in our analysis is to generalize the game to allow the retailer to choose the referral fee as an action. With this, we will build a more accurate and formal model to answer the question of which equilibria are preferable to the retailer. Precisely, we study the following sequential game:
@@
Retailer’s cost $c_{r}$, seller’s cost $c_{s}$
@@
    The retailer and the independent seller play the shared-revenue Bertrand game with parameters $(\alpha,c_{r},c_{s})$
Examining Figure 3, we can think of fixing a cost on the horizontal axis and taking a slice, for instance drawing a vertical line corresponding to $c_{r}=0.9$ on the left plot. With this, the retailer now gets to choose whatever value of $\alpha$ they’d like, which is a point along the vertical line $c_{r}=0.9$, effectively getting to choose which type of equilibrium they would like to observe. We illustrate such a slice in Figure 4 by coloring the background of the plot according to which equilibrium in Figure 3 is observed. As one might expect, we will find that the retailer prefers the red/gray equilibria to the green/blue equilibria. However, despite potentially being a win-win-win, we will find that the retailer does not always choose the red equilibrium when it exists because they may be able to “win more” in the gray region at the expense of customers and/or the independent seller.
@@ 5.1. Refined equilibria of fee optimization game
We will impose the further refinement of subgame perfection (Rasmusen, 1989), which is equivalent to backward induction in this game of perfect, complete information. In order to achieve this, the retailer and independent seller must choose a price configuration that is an admissible, Pareto optimal equilibrium of the shared-revenue Bertrand subgame. However, there are many choices of such price configurations. We will thus define an equilibrium strategy profile $\rho$ for the subgame. Formally, we require that $\rho$ is a strategy profile with the property that for every possible $\alpha,c_{r},c_{s}$, $\rho(\alpha,c_{r},c_{s})\in\mathcal{S}_{p}$ is an admissible, Pareto optimal equilibrium price configuration of the shared-revenue Bertrand game as outlined in (70), (71), and (72) [^5] This function $\rho$ is necessary for subgame perfection and could come from anywhere. Intuitively, we could think of the retailer and independent seller agreeing in advance on what prices they will set for each choice of $\alpha$, given common knowledge of their prices $c_{r},c_{s}$. We could also think of the retailer as more powerful and give them the ability to choose $\rho$, requiring that the independent seller agree to choose the prices outlined in $\rho$ as a condition of the revenue sharing program..
@@
– an admissible, Pareto optimal equilibrium strategy profile of the shared revenue Bertrand game for every set of parameters
@@ 5.2. Interpreting equilibria
To complete our analysis, our natural next step is to find $\alpha_{*}$. We’ll denote as $\pi_{r}^{\text{(eq)}}(\rho(\alpha),\alpha,c_{r},c_{s})$ the payoff function in equilibrium of the shared-revenue Bertrand game for the retailer with strategy $\rho$ and $\pi_{s}^{\text{(eq)}}$ the analogous function for the independent seller. With this notation, it’s clear that
@@
$\alpha_{*}(c_{r},c_{s},\underline{\rho})=\alpha_{r^{*}}\equiv\underline{\alpha}$
In other words, if the retailer and independent seller agree to play strategy $\underline{\rho}$, the retailer will always choose referral fee $\underline{\alpha}=\alpha_{r^{*}}$ in equilibrium of the Fee Optimization Game.
@@
(29) $\alpha_{*}(c_{r},c_{s},\overline{\rho})=\begin{cases}\arg\max_{\alpha}\pi_{r,s}(p_{s^{*}},\alpha)\equiv\overline{\alpha}&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \alpha^{\dagger}_{r}&\text{else}\end{cases}$
@@
To illustrate $\underline{\alpha}$ and $\overline{\alpha}$, we plot typical payoff curves in Figure 4. The background of the plot corresponds to the equilibria in Figure 3. In the left plot, there does exist an $\alpha$ such that $\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})$, so $\alpha_{*}(\overline{\rho})=\overline{\alpha}$. In the right plot, for all $\alpha$ we have $\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$, so $\alpha_{*}(\overline{\rho})=\alpha^{\dagger}_{r}$.
@@ 5.3. General equilibrium strategy profiles
(30) $\pi_{r}\in\begin{cases}\left[\pi_{r,s}(p_{s^{*}},\overline{\alpha}),\pi_{r,s}\left(p_{r^{*}},\underline{\alpha}\right)\right]&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \left[\pi_{r,r}(p_{r^{*}}),\pi_{r,s}\left(p_{r^{*}},\underline{\alpha}\right)\right]&\text{else}\end{cases}$
@@
In light of this general result, our analysis from Section 5.2 has far-reaching implications. Importantly, examining Figure 4, we see that the retailer will always induce an equilibrium where the independent seller fulfills demand at some $p\in\left[\max\left\{p_{s_{\text{ind}}},p_{r^{*}}\right\},\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ depending on $\rho$ (red or gray regions). Interestingly, the socially suboptimal green region where the independent seller fulfills demand at $p_{r_{\text{ind}}}$ is never observed in equilibrium of the Fee Optimization Game.
We do also find that the retailer never sets the fee high enough such that they fulfill demand themselves at price $p_{r^{*}}$ (blue region). However, this is less interesting, because it is an artifact of the relative Pareto optimality criteria that we applied when refining equilibria in Section 4.2; we argued there that the retailer would prefer for the independent seller to fulfill demand at $p^{\dagger}$ rather fulfilling demand themselves at $p_{r^{*}}$.
@@ 6. An outside option for the independent seller
In the Fee Optimization Game, the retailer was free to choose any fee they’d like, and the independent seller would stay in the revenue sharing program irrespective of the fee chosen. However, to make the game more realistic, we should model the independent seller’s choice to enter the revenue sharing program as a negotiation with the retailer. Indeed, in real scenarios the independent seller often has a compelling alternative (outside option) – selling their product independently rather than participating in the revenue sharing program. Formally, we have the following game:
Outside Option Game
Parameters
Retailer’s cost $c_{r}$, seller’s cost $c_{s}$, outside option cost differential $\delta\in(0,c_{r}-c_{s})$
Gameplay
  1. (1)
    The retailer chooses a referral fee $\alpha\in(0,1)$
  2. (2)
    The independent seller chooses one of two options:
    Stay in the retailer’s revenue sharing program
    Leave the retailer’s revenue sharing program and sell on their own
  3. (3)
    The retailer and seller choose their prices:
    If the independent seller chose to stay, they play the shared-revenue Bertrand game
    If the independent seller chose to leave, they play a to-be-defined “leaving” subgame
@@ 6.1. Leaving subgame
To finish defining the Outside Option Game, we must outline the structure of this “leaving” subgame. When not clear from context, we will use superscripts $\cdot^{(\ell)}$ to denote relevant quantities in the leaving subgame and $\cdot^{(o)}$ to denote relevant quantities in the outside option game as a whole. There are many reasonable specifications for this leaving subgame, and more sophisticated models may better capture reality, but as a first step we choose a very simple model: if the independent seller chooses to leave, they will play a standard Bertrand game against the retailer.
@@
Whatever it represents, we will assume that there exists a $\delta>0$ such that the independent seller does not always prefer to leave. We now have a standard Bertrand game with potentially asymmetric costs depending on the value of $\delta$, for which the equilibria have been derived in (Blume, 2003; Kartik, 2011) for instance. However, if $\delta\geq c_{r}-c_{s}$ such that $c_{s}+\delta\geq c_{r}$, the independent seller achieves an equilibrium payoff of 00 if they exercise their outside option, making leaving the revenue sharing program a non-credible threat. This can be seen by adapting the derivation in Appendix G to one where $c_{r}\leq c_{s}+\delta$ instead of $c_{s}+\delta<c_{r}$. Thus, with the constraint that $\delta\in(0,c_{r}-c_{s})$, we show in Appendix G that the admissible equilibrium outcomes of the leaving subgame are
$(p,x)=\left(\min\left\{c_{r},p_{s^{*}}^{(\ell)}\right\},s\right)$
where $p_{s^{*}}^{(\ell)}\equiv\frac{1}{2}(1+c_{s}+\delta)$ is the maximizer of the independent seller’s payoff function in the leaving subgame. Importantly, because the retailer never fulfills demand in equilibrium and there is no revenue sharing, they always achieve an equilibrium payoff of 00 if the independent seller chooses to leave. Because the retailer achieves nonzero payoff in equilibrium otherwise, they would always prefer to choose an $\alpha$ such that the independent seller stays over an $\alpha$ such that the independent seller leaves. Furthermore, notice that the leaving subgame itself is entirely independent of $\alpha$.
6.2. Refined equilibria of outside option game
The analysis in Appendix F.1 about refinements for the Fee Optimization Game equilibria applies to the Outside Option Game equilibria: a strategy of the sequential game is admissible if $\rho$ is admissible, and Pareto optimality implies that the retailer will choose the lowest fee possible to achieve their maximum payoff. Furthermore, because the retailer achieves nonzero payoff if the independent seller chooses to stay and 00 payoff if they leave, Pareto optimality also tells us that if the independent seller is indifferent between staying and leaving, they will choose to stay. To summarize, the admissible, Pareto optimal, subgame perfect equilibrium strategies are:
Equilibrium of Outside Option Game

### v2

@@ 3.4. Summary of Nash equilibria
Equilibrium Price Configurations of Staying Subgame
@@ 4.2. Relative Pareto optimality
$(p_{r},p_{s})\in\left(p_{r^{*}},\left[p^{\dagger},p_{s^{*}}\right]\right)\cup\left(p_{s},\left[p_{s_{\text{ind}}},p^{\dagger}\right]\right)\ ;$ $\text{if }\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}\text{ and }p^{\dagger}\leq p_{s^{*}}$
The equilibrium where the retailer fulfills demand at $p_{r^{*}}$ is Pareto suboptimal relative to the shared-price equilibrium $p^{\dagger}$ where the independent seller fulfills demand. Indeed, by construction of $p^{\dagger}$, in both cases the retailer achieves the same payoff of $\pi_{r,r}(p_{r^{*}})=\pi_{r,s}(p^{\dagger})$ so they are indifferent. However, because $\alpha_{r^{*}}\leq\alpha\leq\alpha^{\dagger}_{s}$ implies that $p_{s_{\text{ind}}}\leq p^{\dagger}$, the independent seller achieves a positive payoff when fulfilling demand at $p^{\dagger}$ while they achieve a payoff of 0 when allowing the retailer to fulfill demand at $p_{r^{*}}$. Thus, they strictly prefer fulfilling demand at $p^{\dagger}$ rather than allowing the retailer to fulfill demand at $p_{r^{*}}$.
As a remark, we note that a reasonable argument could be made for preferring the equilibrium where the retailer fulfills demand at $p_{r^{*}}$. In this case $p_{r^{*}}<p^{\dagger}$, so despite the Pareto optimality argument, the customer will be able to buy the product at a lower price if the retailer fulfills demand at $p_{r^{*}}$ instead of the independent seller fulfilling demand at $p^{\dagger}$. With this lowest price criterion, $p_{r^{*}}$ would in fact be the most likely equilibrium in this case, as $p_{r^{*}}<p_{s_{\text{ind}}}$ as well. Whatever we choose does not substantially affect the results, and the relative Pareto optimality criterion will continue to be sensible in the larger shared-revenue Bertrand game, thus we proceed with this assumption.
@@ 4.3. Equilibrium outcomes
Equilibrium Outcomes of Staying Subgame
@@
A quick glance at (22) and (23) shows that when $\alpha$ is sufficiently high, the independent seller is “priced out” of the market and the retailer will fulfill demand at their single agent price as if the independent seller did not exist. When $\alpha$ is small, the independent seller sells at the retailer’s indifference price, illustrating that they have to prioritize preventing the retailer from undercutting them over maximizing their own payoff. However, when $\alpha$ is in a certain sweet spot, the independent seller may be able to sell at their optimal price in equilibrium. We will see soon that everybody benefits in this regime – if the retailer gets to choose the referral fee, they would prefer to choose a referral fee in this sweet spot rather than one that is too high or too low.
@@ 4.4. Fee and cost structure analysis
4.4.1. $(p_{r^{*}},r)$ - The retailer fulfills demand at its single agent price
@@
4.4.3. $(p_{r_{\text{ind}}},s)$ - The independent seller fulfills demand at the retailer’s indifference price, $p_{r_{\text{ind}}}$
@@
4.4.4. $(p_{s^{*}},s)$ - The independent seller fulfills demand at their single agent price, $p_{s^{*}}$
When $\alpha\in\left[\alpha_{s^{*}},\alpha_{s^{*},r^{*}}\right]$, we have an equilibrium in which the independent seller fulfills demand at its optimal price $p_{s^{*}}$ (red region). In this regime, we always have $p_{s^{*}}\leq p_{r^{*}}$, thus the customers are able to purchase at a lower price than without the revenue sharing program. The independent seller prefers this equilibrium because they are able to sell at their optimal price. Finally, there exist some $\alpha$ in the red region such that the retailer achieves a payoff higher than their optimal single agent payoff, a result which benefits the customers, independent seller, and retailer.
@@ 5. Optimizing the referral fee
In Sections 4.4.2 and 4.4.4, we identified two equilibrium regions where the retailer’s payoff may increase relative to their optimal single agent payoff (without the revenue sharking program). In this section, we allow the retailer to choose the referral fee before playing the staying subgame, which is equivalent to determining which equilibrium is most preferable to the retailer for a given cost configuration. Precisely, we study the following sequential game:
@@
Demand curve $q(p)$, retailer’s cost $c_{r}$, seller’s cost $c_{s}$
@@
    The retailer and the independent seller play the staying subgame with parameters $(q,\alpha,c_{r},c_{s})$
Examining Figure 3, we can think of fixing a cost on the horizontal axis and taking a slice, for instance drawing a vertical line corresponding to $c_{r}=0.9$ on the left plot. With this, the retailer now gets to choose whatever value of $\alpha$ they’d like, which is a point along the vertical line $c_{r}=0.9$, effectively getting to choose which type of equilibrium they would like to observe. We illustrate such a slice in Figure 4 by coloring the background of the plot according to which equilibrium in Figure 3 is observed.
@@ 5.1. Refined equilibria of fee optimization game
We will impose the further refinement of subgame perfection (Rasmusen, 1989), which is equivalent to backward induction in this game of perfect, complete information. In order to achieve this, the retailer and independent seller must choose a price configuration that is an admissible, Pareto optimal equilibrium of the staying subgame. However, there are many choices of such price configurations. We will thus define an equilibrium strategy profile $\rho$ for the subgame. Formally, we require that $\rho$ is a strategy profile with the property that for every possible $\alpha,c_{r},c_{s}$, $\rho(\alpha,c_{r},c_{s})\in\mathcal{S}_{p}$ is an admissible, Pareto optimal equilibrium price configuration of the staying subgame as outlined in (52), (53), and (54) [^7] This function $\rho$ is necessary for subgame perfection and could come from anywhere. Intuitively, we could think of the retailer and independent seller agreeing in advance on what prices they will set for each choice of $\alpha$, given common knowledge of their prices $c_{r},c_{s}$. We could also think of the retailer as more powerful and give them the ability to choose $\rho$, requiring that the independent seller agree to choose the prices outlined in $\rho$ as a condition of the revenue sharing program..
@@
– an admissible, Pareto optimal equilibrium strategy profile of the staying subgame for every set of parameters
@@ 5.2. Interpreting equilibria
To complete our analysis, our natural next step is to find $\alpha_{*}$. We’ll denote as $\pi_{r}^{\text{(eq)}}(\rho(\alpha),\alpha,c_{r},c_{s})$ the payoff function in equilibrium of the staying subgame for the retailer with strategy $\rho$ and $\pi_{s}^{\text{(eq)}}$ the analogous function for the independent seller. With this notation, it’s clear that
@@
$\alpha_{*}(c_{r},c_{s},\underline{\rho})=\alpha_{r^{*}}$
In other words, if the retailer and independent seller agree to play strategy $\underline{\rho}$, the retailer will always choose referral fee $\alpha_{r^{*}}$ in equilibrium of the fee optimization game.
@@
(25) $\alpha_{*}(c_{r},c_{s},\overline{\rho})=\begin{cases}\min_{\alpha}[\arg\max_{\alpha}\pi_{r,s}(p_{s^{*}},\alpha)]\equiv\overline{\alpha}&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \alpha^{\dagger}_{r}&\text{else}\end{cases}$
@@
To illustrate $\alpha_{*}$, we plot typical payoff curves in Figure 4. The background of the plot corresponds to the equilibria in Figure 3. In the left plot, there does exist an $\alpha$ such that $\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})$, so $\alpha_{*}(\overline{\rho})=\overline{\alpha}$. In the right plot, for all $\alpha$ we have $\pi_{r,s}(p_{s^{*}},\alpha)\leq\pi_{r,r}(p_{r^{*}})$, so $\alpha_{*}(\overline{\rho})=\alpha^{\dagger}_{r}$.
@@ 5.3. General equilibrium strategy profiles
(26) $\pi_{r}\in\begin{cases}\left[\pi_{r,s}(p_{s^{*}},\overline{\alpha}),\pi_{r,s}\left(p_{r^{*}},\alpha_{r^{*}}\right)\right]&\exists\alpha\text{ s.t. }\pi_{r,s}(p_{s^{*}},\alpha)>\pi_{r,r}(p_{r^{*}})\\ \left[\pi_{r,r}(p_{r^{*}}),\pi_{r,s}\left(p_{r^{*}},\alpha_{r^{*}}\right)\right]&\text{else}\end{cases}$
@@
Importantly, examining Figure 4, we see that the retailer will always induce an equilibrium where the independent seller fulfills demand at some $p\in\left[\max\left\{p_{s_{\text{ind}}},p_{r^{*}}\right\},\left\{p_{s^{*}},p^{\dagger}\right\}\right]$ depending on $\rho$ (red or gray regions). Interestingly, the socially suboptimal green region where the independent seller fulfills demand at $p_{r_{\text{ind}}}$ is never observed in equilibrium of the fee optimization game. The retailer’s payoff always increases relative to a world where the revenue sharing program did not exist.
We do also find that the retailer never sets the fee high enough such that they fulfill demand themselves at price $p_{r^{*}}$ in equilibrium (blue region). However, this is less interesting, because it is an artifact of the relative Pareto optimality criterion that we applied when refining equilibria in Section 4.2; we argued there that the retailer would prefer for the independent seller to fulfill demand at $p^{\dagger}$ rather fulfilling demand themselves at $p_{r^{*}}$.
@@ 6. An outside option for the independent seller
In the fee optimization game, the retailer was free to choose any fee they’d like, and the independent seller would stay in the revenue sharing program irrespective of the fee chosen. To complete our analysis of the shared-revenue Bertrand game, we now model the independent seller’s choice to leave the revenue sharing program as a single-step negotiation with the retailer. Indeed, in real scenarios the independent seller often has a compelling alternative (outside option) – selling their product independently rather than participating in the revenue sharing program.
@@ 6.1. Leaving subgame
To finish defining the shared-revenue Bertrand game, we must outline the structure of the leaving subgame. When not clear from context, we will use superscripts $\cdot^{(\ell)}$ to denote relevant quantities in the leaving subgame and $\cdot^{(o)}$ to denote relevant quantities in the shared-revenue Bertrand game with the outside option. There are many reasonable specifications for this leaving subgame (Shubik and Levitan, 1980; Tirole, 1988), and more sophisticated models may better capture reality, but as a first step we choose a very simple model: if the independent seller chooses to leave, they will play a standard Bertrand game against the retailer (Bertrand, 1883).
@@
Whatever it represents, we will assume that there exists a $\delta>0$ such that the independent seller does not always prefer to leave. We now have a standard Bertrand game with potentially asymmetric costs depending on the value of $\delta$, for which the equilibria have been derived in (Blume, 2003; Kartik, 2011) for instance. However, if $\delta\geq c_{r}-c_{s}$ such that $c_{s}+\delta\geq c_{r}$, the independent seller achieves an equilibrium payoff of 0 if they exercise their outside option, making leaving the revenue sharing program a non-credible threat. This can be seen by considering the equilibrium in the regime where $c_{r}\leq c_{s}+\delta$ instead of $c_{s}+\delta<c_{r}$. Thus, with the constraint that $\delta\in(0,c_{r}-c_{s})$, the admissible equilibrium outcomes of the leaving subgame are
$(p,x)=\left(\min\left\{c_{r},p_{s^{*}}^{(\ell)}(\delta)\right\},s\right)$
where $p_{s^{*}}^{(\ell)}(\delta)$ is the maximizer of the independent seller’s payoff function in the leaving subgame. Importantly, because the retailer never fulfills demand in equilibrium and there is no revenue sharing, they always achieve an equilibrium payoff of 0 if the independent seller chooses to leave. Because the retailer achieves nonzero payoff in equilibrium otherwise, they would always prefer to choose an $\alpha$ such that the independent seller stays over an $\alpha$ such that the independent seller leaves. Furthermore, notice that the leaving subgame itself is entirely independent of $\alpha$.
6.2. Refined equilibria of shared-revenue Bertand game
The analysis in Appendix F.1 about refinements for the fee optimization game equilibria applies to the shared-revenue Bertrand game with outside option: a strategy of the sequential game is admissible if $\rho$ is admissible, and Pareto optimality implies that the retailer will choose the lowest fee possible to achieve their maximum payoff. Furthermore, because the retailer achieves nonzero payoff if the independent seller chooses to stay and 0 payoff if they leave, Pareto optimality also tells us that if the independent seller is indifferent between staying and leaving, they will choose to stay. To summarize, the admissible, Pareto optimal, subgame perfect equilibrium strategies are:
Equilibrium of Shared-Revenue Bertrand Game